OpenStudy (anonymous):

Solve. x^2 = 10x – 20

OpenStudy (anonymous):

x2-10x+25=20+25 x2-10x+25=45 (x-5)2 = 45 x-5 = square root of 45 answer:5 plus and minus square root of 45

OpenStudy (turingtest):

set equal to zero: x^2-10x+20=0 again, this cannot be factored so you must use the quadratic formula. can you take it from here

OpenStudy (turingtest):

?